Find a partial differential equation by eliminating a and b from the equations of

z = ax + (1 − a)y + b.



1
Expert's answer
2022-03-07T17:08:04-0500

Given "z = ax + (1 \u2212 a)y + b."


Differentiating partially with respect to 'x', "p = \dfrac{\partial z}{\partial x} = a"


Differentiating partially with respect to 'y', "q = \dfrac{\partial z}{\partial y} = 1-a"


Eliminating 'a' from these we get, "p + q = 1", which is the required partial differential equation.